# link prediction; graph neural network; community detection; network analysis; citation graph; deep learning

### 04 Jan 2024 v1

@@ Abstract
This study introduces an innovative approach that integrates community detection algorithms with Graph Neural Network (GNN) models to enhance link prediction in scientific literature networks. We specifically focus on the utilization of the Louvain community detection algorithm to uncover latent community structures within these networks, which are then incorporated into GNN architectures to predict potential links. Our methodology demonstrates the importance of understanding community dynamics in complex networks and leverages the strengths of both community detection and GNNs to improve predictive accuracy. Through extensive experiments on bipartite graphs representing scientific collaborations and citations, our approach not only highlights the synergy between community detection and GNNs but also addresses some of the prevalent challenges in link prediction, such as scalability and resolution limits. The results suggest that incorporating community-level information can significantly enhance the performance of GNNs in link prediction tasks. This work contributes to the evolving field of network science by offering a novel perspective on integrating advanced machine learning techniques with traditional network analysis methods to better understand and predict the intricate patterns of scientific collaborations.
@@ 1 Introduction
A graph is a special structure of data in computer storage that consists of a collection of vertices and a collection of edges. Link prediction is the process of predicting whether there is a possibility of an edge between vertices that are not directly connected to each other in a specific graph structure, and the corresponding weight of that edge Liben2007LinkPrediction . Traditional link prediction methods include Markov chain and machine learning, and have been used for website browsing prediction and literature citation prediction, etc Sarukkai2000MarkovChains Popescul2003StatisticalRelational . However, the two traditional link prediction methods mentioned above are overly dependent on node attributes in the network when modeling, and node attributes in the actual application scenarios are difficult to obtain and cannot be guaranteed to be reliable, so the traditional methods do not have a satisfactory prediction accuracy. Graph embedding techniques offer significant advantages, particularly in their ability to incorporate complex network structures and node relationships more effectively, leading to improved prediction accuracy in various applications Therefore, link prediction algorithms based on graph embedding have gradually become a hot research topic in recent years.
In this paper, we introduce a novel approach that integrates community detection algorithms with Graph Neural Networks (GNN) to enhance the efficiency and accuracy of link prediction in scientific literature networks. Our method utilizes the Louvain algorithm for community detection, which aids in identifying latent community structures within these networks. These structures are then leveraged by GNN models, specifically designed for this task, to predict potential links with higher precision. This integration not only taps into the community dynamics inherent in scientific collaborations but also harnesses the power of GNNs to interpret and utilize these complex network patterns.
The overall approach, succinctly encapsulated in Figure 1, represents a significant step forward in predictive network analysis, particularly in the context of scientific literature where understanding and predicting collaboration patterns can yield valuable insights.
@@ 3.1 Data Collection
In this study, we utilized data extracted from the Web of Science database, focusing on literature pertaining to zinc batteries. Our dataset encompasses a comprehensive collection of 10,187 research articles, identified up to November 2023. The majority of these publications, approximately 9,315 papers, were released post-2010, representing over 90 % of the total corpus. Notably, the articles predominantly appear in leading journals within the field, including the Journal of Power Sources, Chemical Engineering Journal, ACS Applied Materials & Interfaces, Journal of the Electrochemical Society, and Journal of Materials Chemistry A, among others.
For the dataset compilation, a specific search query was employed within the Web of Science database. The query was structured as follows: TI=(Zinc) AND TI=(Anode) OR TI=(Zn) AND TI=(Anode) OR TI=(Zn) AND TI=(Batter*) OR TI=(Zinc) AND TI=(Batter*). This search strategy was designed to capture articles where the title (TI) contained the keywords ’Zinc’ or its abbreviation ’Zn’ in conjunction with ’Anode’ or variations of ’Battery’ (indicated by the wildcard ’*’).
@@ 3.2 Train-Test Split
A network graph is created from an edge dataset to facilitate a balanced approach to link prediction. The process begins with the construction of a graph using edge data, symbolizing connections between nodes. Subsequently, the dataset is partitioned into training and test subsets.
To address the prevalent issue of class imbalance in network data, a novel step is introduced where non-existent links are generated. This generation continues until their count matches the number of existing links in the training set, thereby ensuring an equitable representation of both link types. The final training set, enriched with these synthetic non-links, and a comparably prepared test set, serve as the foundation for further model training and evaluation. This methodology emphasizes accurate prediction of both the presence and absence of links in a network, a critical aspect in network analysis. The resulting sizes of the oversampled training set and the test set are 296,574 and 74,144, respectively, demonstrating the significant expansion of the dataset for balanced training.
@@ 4.2.2 Feature Engineering
In the feature engineering stage of this study, a rigorous process is employed to prepare node attributes for network analysis. This involves filling missing values in textual data, one-hot encoding categorical variables, and applying TF-IDF vectorization to text-based attributes. The integration of these methods results in a multi-dimensional feature space represented as a sparse matrix, encompassing 10,187 samples with 2,026 distinct features. This comprehensive dataset, achieved through meticulous preparation and transformation of node characteristics, is instrumental in constructing a robust foundation for the subsequent predictive modeling in network analysis.
@@ 4.3.2 Graph Attention Networks
Figure 2 includes the training loss data for the GAT and GATv2 models, spanning epochs 0 to 300. The GAT losses are marked with triangles and the GATv2 losses with crosses, using colors consistent with the previous plots.
@@ 4.3.3 Graph Convolutional Networks
Figure 3 illustrates the training loss trajectories of two graph neural network models, GCN and GCNv2, over a series of 300 epochs. The GCN model exhibits a gradual decrease in loss with minor fluctuations, while the GCNv2 model shows a more pronounced decrease, particularly in the early epochs.
@@ 4.3.4 GraphSAGE
Figure 4 includes the updated training loss data for the GraphSAGE model, spanning epochs 0 to 300.
@@ 4.3.5 Graph Neural Networks with Community Detection
In this phase of the study, community detection is integrated into the preprocessing of graph data for enhanced network analysis. The Louvain method, a popular community detection algorithm, is applied to the network graph to identify communities or clusters within it. This method is advantageous for uncovering the modular structure of the network, revealing groups of nodes that are more densely connected internally than with the rest of the network.
After detecting communities, each node is assigned a community label, which is then transformed into a feature using one-hot encoding. These community features are combined with the existing node features, previously processed through methods like TF-IDF vectorization and one-hot encoding for textual and categorical data, respectively. The resulting combined feature set is more comprehensive, now including both individual node characteristics and their community affiliations.
The enriched dataset, now encapsulating both intrinsic node characteristics and their community contexts, provides a robust foundation for GNN models to analyze network structure and predict links more effectively. This advanced preprocessing approach, by integrating community detection, enhances the model’s capacity to understand and leverage the complex, multi-faceted nature of network data.
Figure  5 offers a comprehensive comparison of training loss trajectories across five graph neural network models—GCN, GCNv2, GAT, GATv2, and GraphSAGE, each enhanced with the Louvain method for community detection, over 300 epochs. The initial loss values vary across models, reflecting different starting efficiencies. Notably, GATv2 exhibits significant early fluctuations, peaking at epoch 10. Despite these variations, all models demonstrate a general trend of decreasing loss, indicating progressive learning and improvement. The use of a sequential color palette and distinct markers (circles for GCN, squares for GCNv2, triangles for GAT, crosses for GATv2, and diamonds for GraphSAGE) aids in differentiating each model’s performance. As epochs advance, the models’ loss values begin to converge, suggesting stabilization in learning. This visualization serves as an insightful tool for analyzing the learning behavior and efficiency of different graph neural network models over an extended training period.
@@ 5 Discussion
5.1 Conclusion
In summary, the integration of Louvain community detection into GNN models consistently enhances their performance in link prediction tasks. The improvements across various metrics underline the significance of incorporating community structure into the analysis, providing a more holistic understanding of the network’s underlying patterns. This approach exemplifies the potential of combining node-specific features with broader network structures to enhance the predictive capabilities of GNN models.
@@ 5.3 Future Work
Future research efforts should focus on further refining and advancing the integration of community detection methods with GNN architectures for link prediction. This includes exploring alternative community detection algorithms that may address the limitations of current methods, such as the resolution limit issue in the Louvain algorithm.
Additionally, enhancing the GNN models to better capture and utilize the nuanced structural information provided by community labels could lead to more accurate and efficient link prediction in complex networks. The exploration of dynamic community detection methods to accommodate evolving network structures, as well as the application of these techniques to a wider range of datasets, would also be beneficial.

### 19 Jan 2024 v2

@@ Abstract
This study presents a novel approach that synergizes community detection algorithms with various Graph Neural Network (GNN) models to bolster link prediction in scientific literature networks. By integrating the Louvain community detection algorithm into our GNN frameworks, we consistently enhance performance across all models tested. For example, integrating Louvain with the GAT model resulted in an AUC score increase from 0.777 to 0.823, exemplifying the typical improvements observed. Similar gains are noted when Louvain is paired with other GNN architectures, confirming the robustness and effectiveness of incorporating community-level insights. This consistent uplift in performance—reflected in our extensive experimentation on bipartite graphs of scientific collaborations and citations—highlights the synergistic potential of combining community detection with GNNs to overcome common link prediction challenges such as scalability and resolution limits. Our findings advocate for the integration of community structures as a significant step forward in the predictive accuracy of network science models, offering a comprehensive understanding of scientific collaboration patterns through the lens of advanced machine learning techniques.
@@ 1 Introduction
Graphs represent intricate data structures comprised of vertices and edges, where link prediction aims to foresee potential connections and their strengths within a graph’s structure Liben2007LinkPrediction . While traditional methods like Markov chains and machine learning have found utility in applications such as website browsing and citation prediction Sarukkai2000MarkovChains Popescul2003StatisticalRelational , they often rely heavily on node attributes, which can be challenging to obtain and validate in real-world settings. Consequently, these methods may not always yield high accuracy in predictions. Graph embedding techniques, by contrast, can effectively encapsulate complex structures and relational dynamics, paving the way for more accurate applications of link prediction algorithms, now an active area of research.
In this paper, we introduce a pioneering approach that unifies community detection algorithms with Graph Neural Networks (GNN) to refine link prediction within scientific literature networks. Employing the Louvain algorithm, we reveal latent community structures that GNN models can exploit to enhance link prediction precision. This methodology is particularly potent in scientific collaboration networks, where understanding and forecasting collaborative patterns is crucial. For instance, our approach could significantly improve paper recommendation systems, aiding researchers in finding relevant studies and potential citations when preparing their scientific manuscripts farber2020citation . Such systems are invaluable given the recent exponential growth in scientific publications and the critical role of citing pertinent work.
Our integrated approach, succinctly depicted in Figure 1, marks a substantial advancement in predictive network analysis. It is especially pertinent to the scientific literature domain, where discerning and anticipating collaboration and citation dynamics can offer profound insights, potentially fostering novel research intersections and driving the progress of scientific inquiry forward.
@@
The structure of the paper is as follows: Section 2 discusses related work to frame the current advancements in link prediction and graph neural networks. Section 3 explains the datasets, data collection, and preprocessing steps. Section 4 delves into our methods, firstly examining heuristic-based link prediction techniques, then exploring machine learning methods, followed by a detailed look at our advanced graph neural network approaches with community detection. Section 5 discusses the broader implications, limitations, and avenues for future research arising from our study. The paper concludes in section 6, summarizing our contributions and reflecting on the potential impact on network analysis in the realm of scientific literature.
@@ 2.1 Graph Neural Network
Building on the extensive survey of graph embedding techniques presented in Section 2, our study specifically focuses on the advancements of GNN methods, particularly GAT, GCN, and GraphSAGE. These methods represent the most directly related prior work to our study due to their state-of-the-art performance in various graph-based tasks and their relevance to our proposed approach. In the context of these foundational GNN models, our work introduces a novel enhancement by integrating the Louvain method for community detection. This integration significantly boosts the models’ performance across multiple metrics, including precision, recall, F1 score, and AUC score. The Louvain method’s ability to uncover latent community structures provides an additional layer of information, enriching the original node features and thereby enabling the GNN models to achieve a deeper understanding of the graph’s topology. This enhancement aligns with the recent trends in deep learning that prioritize the incorporation of structural and contextual information within the learning process, affirming our contribution to the evolving landscape of graph neural network research.
@@ 2.2 Link Prediction
In addressing the evolving landscape of link prediction, our study employs the same rigorous evaluation metrics as previous research, to ensure comparability and consistency. We have established a robust baseline by encompassing heuristic-based methods and machine learning approaches, which sets the stage for our advanced GNN models.
@@ 3.1 Data Collection
In this study, we harnessed the Web of Science database to construct a dataset central to the technological advances in zinc batteries, an area increasingly vital due to the surge in demand for green energy solutions and new energy electric vehicles. Recognizing zinc’s relative abundance compared to lithium, and its potential for industrial application, our search targeted a corpus reflective of the most recent and impactful research.
We retrieved a comprehensive collection of 10,187 research articles using a refined search strategy based on relevant keywords that signify the core technologies and innovations in the field: TI=(Zinc) AND TI=(Anode) OR TI=(Zn) AND TI=(Anode) OR TI=(Zn) AND TI=(Batter*) OR TI=(Zinc) AND TI=(Batter*) shi2020overview . This strategy was meticulously designed to capture the breadth and depth of contemporary zinc battery research, which is predominantly represented in publications post-2010, accounting for over 90% of the total corpus and indicating a burgeoning interest and development in this domain. The dataset includes articles until November 2023, with a significant concentration of papers from authoritative journals such as the Journal of Power Sources, Chemical Engineering Journal, ACS Applied Materials & Interfaces, Journal of the Electrochemical Society, and Journal of Materials Chemistry A. This approach ensures that our analysis is grounded in the most current and influential studies, providing a robust foundation for our subsequent predictive modeling and network analysis.
@@ 3.2 Train-Test Split
In our study, a network graph was generated from an edge dataset, representing the connections between nodes. This graph was then divided into training and test sets, adhering to a standard 80:20 split ratio, resulting in a training set comprising 80% of the data and a test set with the remaining 20%.
To counter the issue of class imbalance typically present in network data, we employed a novel approach to generate synthetic non-links. This procedure was continued until the number of non-links equaled the number of actual links within the training set, ensuring a balanced distribution of classes. This method was mirrored in the preparation of the test set. As a result, the training set expanded to include 296,574 samples, while the test set contained 74,144 samples. This deliberate oversampling strategy for the training set was critical for developing a model that can accurately predict both the existence and absence of links, which is essential for robust network analysis.
@@ 4.2.2 Feature Engineering
The feature engineering phase in our study is a pivotal step towards enabling effective network analysis. Initially, we standardize numerical data using a standard scaler to ensure mean normalization and variance scaling. For categorical variables, a one-hot encoding scheme harris2012digital is applied to convert these variables into a binary vector representation, avoiding any implicit ordering that could be mistakenly inferred by the models. Textual data undergoes a transformation via TF-IDF (Term Frequency-Inverse Document Frequency) vectorization sparckjones1972statistical , a technique that reflects the importance of words relative to a document collection, hence allowing for the nuanced representation of text features.
The culmination of these preprocessing techniques is a multi-dimensional feature space, structured as a sparse matrix to manage the high dimensionality efficiently. The final feature space encompasses a total of 10,187 samples, each characterized by a feature vector consisting of 2,026 distinct elements. This feature space’s dimensionality was carefully curated to capture the rich, underlying structures of the node attributes while avoiding the pitfalls of overfitting that can occur with excessively high-dimensional data. The resulting dataset, formed through careful curation and transformation of node characteristics, provides a comprehensive and robust foundation for the subsequent phases of predictive modeling in our network analysis.
@@ 4.3 Graph Neural Networks
Figure 2 illustrates the process by which the GNN models utilize subgraphs observed from the input graph to predict links. This visual representation clarifies the methodology underlying the application of various GNN variants and demonstrates the nuanced approach of our analysis.
Figure 2: The general methodology of different GNN variants.
@@ 4.3.1 Overview
The confusion matrices presented in figure 7 correspond to the models listed in table, with the same ordering. They illustrate the performance of each model on a consistent test set, with a fixed random seed of 42 to ensure reproducibility. The left side of each matrix pertains to models using only the original node features, while the right side shows the performance after the integration of community detection features via the Louvain method.
@@ 4.3.2 Graph Attention Networks
Figure 3 presents the training loss trajectories for both the GAT and GATv2 models over a span of 300 epochs, providing a visual representation of the models’ learning progress. The GAT model’s loss, denoted by triangles, shows an initial rapid decline, stabilizing as the epochs increase, which is indicative of the model effectively learning from the training data. Similarly, the GATv2 model, begins with a higher loss that quickly descends, albeit with some early fluctuations, before it too stabilizes. This pattern is reflective of the GATv2’s advanced attention mechanisms adapting to the graph structure over time. The colors selected for the plot are consistent with those used in previous figures for ease of comparison.
@@
In our experimental validation, the hyperparameters for the GAT and GATv2 model were optimized through comprehensive testing. The final configuration, yielding the most efficacious results, is delineated in table 4 and 5.
@@ 4.3.3 Graph Convolutional Networks
Figure 4 displays the training loss trajectories for the GCN and GCNv2 models across 300 training epochs. The GCN model’s losses are indicated with circles and exhibit a gradual descent with some variability, suggesting an incremental learning process. In contrast, the GCNv2 model, marked with squares, shows a steeper initial decline in loss, indicating a rapid adaptation to the data. This steep decline in the early epochs for GCNv2 can be attributed to the enhanced convolution layers and linear transformation layer that allow it to capture complex graph structures more effectively.
@@
The hyperparameter tuning undertaken for the GCN and GCNv2 model was systematic, aiming to ascertain an optimal set of configurations for enhancing model performance. The resulting parameters, determined to be most effective through empirical testing, are in table 6 and 7.
@@ 4.3.4 GraphSAGE
Figure 5 presents the training loss trajectory for the GraphSAGE model over 300 epochs, providing insight into the model’s learning performance throughout the training phase. The loss curve, depicted with a yellow line, begins with a sharp decline, indicating a rapid initial learning phase, and then transitions into a more gradual descent. This suggests that GraphSAGE quickly assimilates the local graph structures and then refines its understanding of the data over time. The relative stability of the loss after the initial epochs showcases the efficiency of GraphSAGE’s neighborhood sampling and aggregation functions in generating informative node embeddings.
@@
The hyperparameter tuning process was executed with the goal of identifying the most effective model parameters. The selected hyperparameters for the Enhanced GraphSAGE model are detailed in the table 8.
@@ 4.3.5 Graph Neural Networks with Community Detection
In the enhanced methodology of our study, the preprocessing of graph data is elevated through the integration of the Louvain community detection algorithm blondel2008fast . The Louvain method operates on the principle of modularity optimization. Mathematically, it seeks to maximize the modularity $Q$, defined as
$Q=\frac{1}{2m}\sum_{ij}\left[A_{ij}-\frac{k_{i}k_{j}}{2m}\right]\delta(c_{i},c_{j}),$
where $A$ is the adjacency matrix of the network, $m$ is the sum of all edge weights, $k_{i}$ and $k_{j}$ are the degrees of nodes $i$ and $j$, and $\delta$ is the Kronecker delta function that equals 1 if nodes $i$ and $j$ are in the same community and 0 otherwise.
By applying this algorithm, each node is endowed with a community label that encapsulates its modular connectivity within the graph. These labels are then encoded into a one-hot vector, expanding the original node feature set—previously vectorized by TF-IDF and one-hot encoding techniques—by appending this new community structure information. Consequently, the feature space dimensions increase from 2,026 to 3,315, enriching the input data for the GNN models.
The GNN models—GCN, GCNv2, GAT, GATv2, and GraphSAGE—are subsequently trained on this augmented dataset, harnessing not only the intrinsic node features but also the macro-structural properties encoded by the Louvain method. This dual representation fosters a deeper interpretative capacity within the models, leading to an improved learning trajectory as observed in Figure  6. The loss trajectories for models trained with community detection features indicate a more pronounced and rapid stabilization, highlighting the effectiveness of the Louvain method in enhancing the network analysis capabilities of GNNs.
@@ 5.2 Future Work
As we look to the future, the refinement of integrating community detection with GNNs stands as a promising avenue for advancing link prediction techniques. Beyond simply enhancing the accuracy of link predictions, the amalgamation of these two methodologies has the potential to significantly influence the broader field of network analysis. The nuanced understanding of community structures within networks could revolutionize the way we approach complex networked systems, enabling a deeper comprehension of their evolution, resilience, and underlying dynamics.
In the realm of scientific literature, the implications of this research are manifold. By improving paper recommendation systems, our approach can streamline the research process, enabling scholars to discover relevant works with unprecedented precision. This is particularly vital in an era where the sheer volume of publications can overwhelm researchers. Moreover, by anticipating collaboration and citation patterns, we can not only track but also predict the progression of scientific discourse, potentially catalyzing new research intersections and collaborations that might have otherwise remained undiscovered.
The integration of community detection with GNNs could also be extended to other domains where network analysis plays a critical role, such as social network analysis, epidemiology, and bioinformatics. For instance, in social networks, this methodology could enhance the detection of community structures and their changes over time, providing insights into the formation of social groups and the spread of information or misinformation. In epidemiology, it can aid in understanding and predicting the spread of diseases within and between communities, potentially informing public health strategies and interventions. In bioinformatics, it could improve the prediction of protein-protein interactions, which is crucial for understanding cellular processes and the development of drugs.
Furthermore, exploring alternative community detection algorithms can address current limitations and open new research directions, such as the study of community evolution in dynamic networks. The potential to adapt GNN models to better incorporate temporal and structural changes within communities could lead to groundbreaking developments in dynamic network analysis.
In summary, the integration of community detection methods with GNN architectures represents not just a methodological enhancement in link prediction tasks but also a strategic shift towards a more sophisticated understanding of networks. This could lead to developing more intelligent systems that can adapt to and predict changes within networks, offering significant benefits across various scientific and social domains. Hence, our future work aims not only to improve upon the technical aspects of our approach but also to explore its broader applications, thereby contributing to the progressive evolution of network science and its capacity to address complex real-world problems.
6 Conclusion
In conclusion, our study provides compelling quantitative evidence that the integration of the Louvain community detection method with various GNN models results in robust performance enhancements in link prediction tasks. For instance, when Louvain is applied to GNN models, we observe significant improvements in AUC scores, such as an increase from 0.777 to 0.823 for GAT and from 0.892 to 0.917 for GCNv2. These improvements are consistent across different GNN architectures, underlining the efficacy of incorporating community structures into predictive modeling.
Furthermore, our approach of Louvain-enhanced GNNs not only consistently outperforms traditional machine learning methods, as evidenced by the highest AUC score of 0.917 achieved by GCNv2 with Louvain, but also demonstrates superior results compared to heuristic-based methods, with the Resource Allocation index achieving an AUC score of 0.847. Such findings advocate for a paradigm shift towards the adoption of community-aware GNN models for link prediction, highlighting the importance of a synergistic perspective that blends node-specific features with macroscopic network structures. This study, therefore, concludes that leveraging the latent community information within GNN architectures can substantially elevate the performance of link prediction models, setting a new benchmark for future research in this domain.
@@ 6 Conclusion
ReLU Rectified linear unit
ELU Exponential linear unit
@@
[…]
Appendix A Detailed Architecture and Hyperparameter Tuning of Enhanced GNN Models
This appendix provides a detailed overview of the architecture of our Enhanced GNN models, as well as the hyperparameter tuning process that was undertaken to optimize their performance. These details are crucial for the reproducibility of our research and for providing insights into the potential of deep learning to enhance network analysis.
A.1 Graph Attention Networks
The architecture of our Enhanced GAT model is meticulously constructed to leverage the attention mechanism for node feature refinement. Beginning with the input layer, the model accepts node features and advances them through a graph attentional layer (conv1), which employs four attention heads to project the input into a higher dimensional space, here chosen as 64. Each head computes separate features which are then concatenated, allowing the model to attend to information from different representation subspaces. This layer is followed by an Exponential Linear Unit (ELU) activation function and a dropout layer with a rate of 0.6. The second graph attentional layer (conv2) consolidates the features from the previous layer into a 16-dimensional space. In this final layer, a single attention head is employed to focus on the most salient features for the link prediction task. The utilization of dropout before each attentional operation ensures that the model remains robust to overfitting.
Table 4: Optimal Configuration for the Enhanced GAT Model
Parameter Value
Learning Rate
0.01
Dropout Rate
0.6
Number of Attention Heads
4 (conv1), 1 (conv2)
Number of Epochs
300
Optimizer
Adam
Loss Function
Binary Cross-Entropy (BCELoss)
A.2 Graph Attention Networks Version 2 (GATv2)
The Enhanced GATv2 model is an iteration on the GAT architecture, refined to harness the second version of the graph attention mechanism. At its core, the input layer receives the node features and directs them through the first graph attentional layer (conv1), which utilizes four attention heads to expand the feature representation, subsequently concatenated to provide a rich, multidimensional feature space. Each attention head, operating with an individual weight matrix, allows the model to diversify the focus on different neighborhoods of the input graph, thereby enriching the feature extraction process. The conv1 layer outputs are then activated using an Exponential Linear Unit (ELU) to maintain a non-linear transformation while controlling for neuron saturation, followed by a dropout mechanism set at a rate of 0.6 to enhance model generalization.
Continuing the feature refinement, the architecture employs a second graph attentional layer (conv2) that consolidates the previously expanded features into a singular, 48-dimensional space conducive to the link prediction task. This layer uses a single attention head, concentrating the model’s focus on the most pertinent features. As in the first layer, the dropout is applied post-activation to regularize the learning process.
The GATv2 model structure—starting with conv1 followed by ELU and dropout, and progressing through conv2 with the same sequence—presents an advanced approach to leveraging attention mechanisms for graph neural networks. This architecture is particularly adept at capturing and processing complex dependencies within graph data, a critical capability for predictive tasks in network analysis.
Table 5: Optimal Configuration for the Enhanced GATv2 Model
Parameter Value
Learning Rate
0.02
Dropout Rate
0.6
Number of Attention Heads
4 (conv1), 1 (conv2)
Number of Epochs
Determined by validation performance
Optimizer
Adam
Loss Function
Binary Cross-Entropy (BCELoss)
A.3 Graph Convolutional Networks
The Enhanced GCN model’s architecture is composed of a sequence of graph convolutional layers designed to systematically refine node features for optimal representation. The architecture initiates with an input layer that takes node features and subjects them to a graph convolutional operation (conv1), mapping them to a 512-dimensional hidden space. This is immediately succeeded by a ReLU activation function to introduce non-linearity, followed by a dropout operation with a rate of 0.5 to mitigate overfitting. The process proceeds to the second graph convolutional stage (conv2), which further processes the feature vector, reducing its dimensionality to 128. Similar to the previous stage, this is paired with a subsequent ReLU activation and dropout. The architecture is completed by a third graph convolutional layer (conv3), which provides an additional transformation, producing a 32-dimensional feature vector as output. The arrangement of the layers—conv1, ReLU, dropout, conv2, ReLU, dropout, and then conv3—facilitates a comprehensive feature learning mechanism, integral for the Enhanced GCN model’s adeptness at link prediction tasks within network analysis.
Table 6: Optimal Configuration for Our Enhanced GCN Model
Parameter Value
Learning Rate
0.01
Dropout Rate
0.5
Number of Epochs
300
Optimizer
Adam
Loss Function
Binary Cross-Entropy (BCELoss)
A.4 Graph Convolutional Networks Version 2 (GCNv2)
The Enhanced GCNv2 model introduces additional sophistication to the foundational GCN architecture by incorporating linear transformation layers and advanced convolutional operations. The input layer commences the process by receiving node features, which are then subjected to a linear transformation layer to reduce their dimensionality to 48. Following this, a sequence of GCNv2 convolutional layers (GCN2Conv) are applied. These layers are equipped with parameters $alpha$ and $theta$, which are essential in controlling the mixture of the initial node features and the propagated features, as well as the trade-off between the two. Each GCN2Conv layer operates on the principle of iteratively refining node features while maintaining a connection to the original features, thus enabling a deeper level of feature extraction without over-smoothing.
The model’s architecture is characterized by three successive GCN2Conv layers, each followed by a ReLU activation function to inject non-linearity, and a dropout layer with a rate of 0.5, serving as a regularizer. This structured progression of layers—linear transformation, followed by three iterations of GCN2Conv, ReLU, and dropout—provides a potent mechanism for feature learning that captures both local and global graph structures pertinent to link prediction.
Table 7: Optimal Configuration for the Enhanced GCNv2 Model
Parameter Value
Learning Rate
0.02
Dropout Rate
0.5
Alpha
0.3
Theta
0.7
Number of Epochs
Defined by early stopping or maximum epochs
Optimizer
Adam
Loss Function
Binary Cross-Entropy (BCELoss)
A.5 Graph Sample and Aggregate Networks (GraphSAGE)
The Enhanced GraphSAGE model’s architecture is defined by a sequence of graph convolutional layers that are strategically structured to optimize the representation of node features for link prediction. The model begins with the first GraphSAGE convolutional layer (conv1), which takes the input node features and expands them to a 512-dimensional hidden layer. This is followed by a ReLU activation to introduce non-linearity, and a dropout layer with a rate of 0.5, designed to prevent overfitting. The model then proceeds to a second GraphSAGE convolutional layer (conv2) that reduces the dimensionality from 512 to 128, applying the same sequence of ReLU activation and dropout. The final layer (conv3) is an additional GraphSAGE convolutional layer, further condensing the features to a 32-dimensional hidden space. This progression—conv1, ReLU, dropout, followed by conv2, ReLU, dropout, and concluded with conv3—ensures a deep and comprehensive feature learning process, which is critical for capturing the complex patterns inherent in network data for effective link prediction.
Table 8: Optimal Configuration for the Enhanced GraphSAGE Model
Parameter Value
Learning Rate 0.01
Dropout Rate 0.5
Hidden Layer Dimensionality 32
Number of Epochs 300
Optimizer Adam
Loss Function Binary Cross-Entropy (BCELoss)
Appendix B Analysis of Confusion Matrices for GNN Models
The inclusion of confusion matrices in this study serves to provide a detailed assessment of the classification performance of various GNN models, both with and without the enhancement of the Louvain method. Confusion matrices offer a comprehensive visualization of the model’s predictive capabilities, displaying the true positives, false positives, true negatives, and false negatives. This information is crucial for understanding the specific types of errors made by the models and for evaluating their performance beyond aggregated metrics like precision, recall, and F1 score.
Figure 7: Confusion matrices of different models.
Appendix C Experimental Setup and Reproducibility
In the interest of reproducibility and to support other researchers in building upon our findings, we hereby detail the computational resources and frameworks utilized in our experiments. Our models were implemented using PyTorch pytorch2019 , a widely-adopted machine learning framework known for its flexibility and efficiency in research prototyping and production deployment. All experiments were executed on a system outfitted with a single RTX 4060 GPU, ensuring a high degree of computational performance.
[…]